# Time-domain study of coupled collective excitations in quantum materials

### v1

@@ IV.4 Phonon-exciton coupling
When a coherent phonon mode is driven out of equilibrium following an optical pump, the resulting atomic motions can in turn modulate the optical properties in a coherent manner. In materials whose optical response is strongly modified by the presence of excitons — or bound electron-hole pairs — the exciton binding energy 211, 212, 213, lifetime 214, 215, and transport behavior 216 can also be changed. Such phonon-exciton coupling has been widely studied in quantum dots and quantum wells 217, 218, 219, and more recently, in complex bulk and atomically-thin materials 211, 212, 213. By comparing the exciton lineshape in broadband absorption or photoluminescence spectra before and after the photoexcitation event, one can observe how excitons respond to phonons in the time domain, yielding insights for tuning exciton properties via ultrafast phonon engineering.
@@
For example, in anatase TiO2, a bulk semiconductor, Baldini et al. used photoinduced coherent acoustic phonons to control its excitons at room temperature, observing an exciton energy shift up to 30 meV, which is one of the largest reported in solids under a time-dependent perturbation 211. In a prototypical two-dimensional lead iodide perovskite, (PEA)2PbI4 (PEA = phenylethylammonium), it was further discovered that different excitons have varying coupling strengths with different coherent phonons, suggesting distinct polaronic characters of each exciton in this system and possibly in a broader class of hybrid semiconductors 212. In the atomically thin limit, Trovatello et al. also showed a pronounced phonon-exciton coupling effect in a monolayer MoS2, whose excitons are strongly affected by a photoinduced coherent phonon 213. Using transient absorption spectroscopy, the authors studied the interaction between out-of-plane phonon modes and excitons, noting a marked enhancement in the phonon oscillation amplitude at resonant exciton frequencies.
@@
The aforementioned studies of phonon-exciton coupling, however, lack spatial information, which is important for learning the transport behavior of excitons under the influence of phonons. Recently, Peng and coworkers demonstrated microscopic imaging of directional transport of interlayer excitons in bilayer WSe2 with the assistance of surface acoustic waves (ref. 216), and their experimental scheme is summarized in Fig. 9a. In this geometry, excitons were optically excited by a continuous-wave laser on the left edge of the sample and probed at different locations via photoluminescence imaging, giving rise to a spatially-resolved mapping of exciton propagation. As shown in Fig. 9b, the excitons were launched near the position at 20 m, where the strongest photoluminescence intensity appears. In the absence of a surface acoustic wave, the exciton is largely localized (Fig. 9b). However, by launching a surface acoustic wave using an interdigital transducer, excitons can travel a long distance (m, Fig. 9c), which is much longer than their typical diffusion length. Despite the lack of temporal resolution in this measurement, one can imagine a spatiotemporal experimental scheme by using a pulsed laser to launch the excitons. Crucially, this experiment introduces a new way for controlling exciton transport via electrically-excited or photoinduced acoustic waves, a versatile approach that can be adapted to many other 2D semiconductors and heterostructures.
@@ IV.5 Other collective excitation couplings
Despite the special focus on phonons as one of the coupled modes in the preceding sections, the full combinatorics of collective excitation coupling has been applied to many other scenarios such as magnon-magnon 222, 221, 220, 223, exciton-magnon 224, plasmon-polariton 225, 144 among other interactions. In this section, we discuss a few examples that showcase the diversity of collective excitation couplings and their role in determining the ultrafast dynamics of quantum materials.
@@
As an analogy of nonlinear phononics that opens a new pathway for coherent manipulation of the crystalline structure, magnon-magnon coupling provides an avenue for controlling the spin dynamics. Utilizing 2D THz spectroscopy, Blank and coworkers revealed the down-conversion coupling of two magnon modes in the canted antiferromagnet FeBO3 (ref. 222). In another canted antiferromagnet, ErFeO3, Zhang and coworkers also demonstrated a unidirectional up-conversion coupling between distinct magnon modes (ref. 221). The canted spins in these materials result from the competition between antiferromagnetic interaction and Dzyaloshinskii-Moriya interaction, yielding a small net magnetic moment . As a result, there are two primary cooperative motions of the sublattice spins, corresponding to two distinct magnon modes sketched in Fig. 10a: the quasi-ferromagnetic (qFM) mode where precesses around the -axis, and the quasi-antiferromagnetic (qAFM) mode where oscillates. Given the different contributions to from the two modes, they can be selectively excited by tuning the direction of the magnetic field in the incident THz pump pulse, and their dynamics can be detected in the radiated free-induction decay signals. For example, when the magnetic field of the emitted THz pulse is parallel to that of the excitation field (), qFM mode and qAFM mode are selectively excited and observed when and , respectively (upper panel in Fig. 10b). This observation is expected as qFM involves a precession in the - plane while qAFM involves an amplitude oscillation along . Surprisingly, when , qAFM mode was still observed with (lower panel in Fig. 10b), which was not expected if there were only linear interactions between the driving THz magnetic field and magnons. To elicit the nonlinear interactions, 2D THz spectroscopy measurements were employed, where the nonlinear spectrum reveals a unidirectional up-conversion process from qFM to qAFM. As shown in Fig. 10c, the unidirectionality is evidenced by a single peak at the qFM excitation frequency and qAFM detection frequency, and importantly, a peak is absent when the corresponding excitation and detection frequencies are interchanged. Besides magnon up-conversion that echoes the phonon up-conversion seen in SrTiO3 (Fig. 5c, ref. 163), the sum and difference frequencies between these two magnon modes have been demonstrated in a cousin rare-earth ferrite, YFeO3, which are featured by a plethora of off-diagonal peaks in the nonlinear 2D terahertz spectrum in Fig. 10d (ref. 220). Like nonlinear phononics, these nonlinear effects of magnons demonstrated in the time domain help usher in new ways to coherently manipulate the spin degree of freedom in magnetic materials.
@@
The analogy between phonons and magnons goes beyond anharmonic couplings. Like phonons, a magnon wavepacket can propagate in materials, which forms the bedrock for spintronic and magnonic applications. When coupled with other collective excitations, magnon propagation can be modified. For instance, for magnons that are coupled with acoustic phonons, theoretical 226 and experimental 227, 228 works have shown that magnon polarons travel over long distances with a faster group velocity than magnons alone. In another study of a van der Waals antiferromagnetic semiconductor CrSBr, magnons were observed to coherently travel beyond several micrometers with a very fast group velocity, an effect that Bae and coworkers interpreted as the coupling between the magnon and acoustic phonon branches 224, 229. What is interesting about this experiment is that coherent magnons were probed by laser pulses tuned to the exciton resonances in CrSBr, demonstrating an efficient optical access to spintronic information due to strong exciton-magnon couplings in this material. A subsequent experiment suggested that optically launched magnon propagation in CrSBr and likely other van der Waals magnet is instead mediated by long-range magnetic-dipole interactions 230, calling for additional spatiotemporally-resolved studies to understand magnon dynamics in this important class of materials.
@@
Non-local propagation effects mediated by the coupling between collective excitations are not restricted to magnons. Other examples include exciton propagation with the aid of surface acoustic waves in a GaAs quantum well 231, 232 or in a transition metal dichalcogenide 216, and surface-plasmon-polariton propagation at a metal-dielectric interface 233. Similar to a phonon-polariton, a plasmon-polariton is a hybridized excitation between charge oscillations and light. More specifically, when light illuminates a metal-dielectric interface, the oscillating electric field of light drives a collective oscillation of the surface charge, which in turn radiates electromagnetic waves. As a result, a hybridized wave, namely surface-plasmon-polariton that involves both the charge motion on the metal surface and electromagnetic field in the dielectric, propagates along the interface. To observe and control its propagation dynamics, PINEM is one of the leading techniques 146, 144. Figure 11a shows the spatiotemporally-resolved imaging of surface-plasmon-polaritons induced by near-infrared laser pulses on an isolated silver nanowire suspended on a graphene film 144, where surface-plasmon-polaritons are evidenced by the additional peaks that appear in the electron energy loss spectra when the light and electron pulses are temporally overlapped. These peaks result from the quantized gain and loss of energy when electrons interact with the surface-plasmon-polariton field. Using energy-filtered electrons, the temporal dynamics of the surface-plasmon-polariton field was imaged with high spatial resolution as it forms an evolving interferometric standing wave pattern (Fig. 11b–f). Moreover, with different light polarizations, different surface-plasmon-polariton field distributions were obtained, demonstrating the potential to control surface-plasmon-polariton using light.
@@ V Summary and outlook
The study and manipulation of coupled collective excitations by ultrafast light pulses offer a new perspective to disentangle intertwined degrees of freedom that are difficult to isolate in equilibrium. In this review, we have highlighted several couplings that both yield an in-depth understanding of the equilibrium properties and help one to coherently control the nonequilibrium dynamics in the host material.
While examples mentioned so far only include two participating collective excitations, multipartite couplings have been theoretically pondered, such as a hybrid exciton-phonon-photon that gives rise to a “phonoriton” in an optical cavity (Fig. 12a) 236, 234, and experimentally realized, such as the hybrid plasmon-exciton-photon — known as “plexcitons” — in WS2 (Fig. 12b), where cavity effects were achieved through nano-patterning (refs. 235, 237). These studies fall into a larger body of works that aims to manipulate quantum materials via cavity control, in which the light-matter coupling strength can be greatly enhanced by replacing the classical laser fields with quantum-mechanical photon modes. One of the most celebrated examples is microcavity exciton–polaritons in GaAs quantum wells, which arise from the strong coupling between photons and excitons and exhibit remarkable effects such as Bose-Einstein condensation 238, 239, 240. More recently, a multitude of theoretical proposals have been put forward for tailoring the unique properties of quantum materials via cavity design 241, 242, with experimental progress in the study of topological state 243, superconductivity 244, and phase transition 245 in different material platforms.
@@
First, advances in engineering the waveform of the excitation light pulse — such as its polarization, carrier envelope phase, and orbital angular momentum — can greatly expand the tool kits for resonant excitation of collective modes and hence a deterministic control over material properties. A recent example is the synthesis of an intense, circularly-polarized THz pulse used to induce circular motions of atoms, otherwise known as chiral phonons, which can in turn induce a large effective magnetic field 76, 71, 70, 69. In particular, Davies et al. demonstrated that the induced field can significantly alter the magnetic order even at considerable distances from the phonon creation point 70, underscoring the need for a deeper insight into chiral phonon transport. Equipped with the capability to deterministically excite particular phonon and magnon modes using the mechanisms introduced in this review, one can envision future experiments where chiral phonons and magnons are created at distinct sample locations, and their mutual interaction over space and time can be used to gain a mechanistic understanding of the propagation mechanism for realizing ultrafast non-local control of magnetic ordering.
@@
Second, from a technical point of view, great care needs to be taken when comparing the dynamics of different collective excitations probed by distinct experimental setups, making a quantitative study of their couplings challenging. An example of this issue arises when one extracts the electron-phonon coupling strength by comparing the electronic band change and atomic displacement, which are probed by time-resolved photoemission and X-ray diffraction respectively 19, 22. For the same incident pump fluence, the effective excitation fluence can differ substantially because photoemission is a surface-sensitive probe (sub- to few nanometers), whereas X-rays can probe much deeper into the bulk (tens of nanometers to micrometers). Therefore, careful fluence calibrations are needed when comparing these probes unless monolayer samples are used to eliminate the variance. An alternative route is to look for complementary probes with a similar probing depth, such as high-energy electron diffraction and extreme ultraviolet absorption spectroscopy, both operated under a transmission geometry. While the former is sensitive to the lattice degree of freedom, the latter excels in capturing carrier, magnetic, and excitonic excitations with an exceptional time resolution 25, 246. One can envision a chamber design that integrates both probes in a single setup, making it possible to concurrently access lattice and electronic information following the same excitation pulse.
@@
Third, while the review has largely focused on coherent collective dynamics with a well-defined phase, not all light-induced collective excitations are coherent or fully reproducible over multiple pump pulses, particularly in systems with spatial inhomogeneities where defects and disorders are prevalent 247, 248, 249, 250. In this aspect, techniques with both spatial and temporal resolutions that operate in the single-shot mode are important for revealing the local, incoherent, and stochastic dynamics that vary from pulse to pulse, as recently demonstrated in the study of magnetic fluctuations in magnetic multilayer thin films 138, 139 and the heterogeneous insulator-to-metal transition in VO2 (ref. 140), both of which utilized a coherent X-ray source in large-scale user facilities. It is worth noting that stochastic dynamics can also be accessed without relying on single-shot detection 251, which makes it possible to study fluctuating systems with the stroboscopic scheme employed in the vast majority of ultrafast measurements. Along with rapid progress in facility-based ultrashort coherent X-ray sources, we also envision two key developments in table-top experiments for probing local and incoherent collective modes. In the extreme spatial scale, innovations in scanning tunneling microscopy has enabled new ways to image collective excitations at the atomic scale 141, 142, paving the way for controlling these local dynamics through creative engineering of nanoscopic defects. In the extreme temporal scale, ultrafast electron diffraction and microscopy that have witnessed a rapid improvement of their temporal resolution 252, 109, 253, 254 would provide a useful platform to study the initial emergence of incoherent phonons following photoexcitation 255, 256, which play an important role in determining the ensuing defect structure 53 and energy flow in the host material system 257.
@@
As we conclude this review that has touched a diverse set of experimental techniques and observables, it is an apt place to be reminded that the overarching goal of studying coupled collective excitations is to uncover a set of unifying principles that govern the interaction among the different degrees of freedom in solids. While studying collective excitations is certainly not the only way to unravel the intricate microscopic interactions, they offer us a way to describe a many-body system without the need to keep track of individual dynamics of an astronomical number of particles, making it tractable to study and engineer material properties both in and out of equilibrium.

### v2

@@ IV.4 Phonon-exciton coupling
When a coherent phonon mode is driven out of equilibrium following an optical pump, the resulting atomic motions can in turn modulate the optical properties in a coherent manner. In materials whose optical response is strongly modified by the presence of excitons — or bound electron-hole pairs — the exciton binding energy 215, 216, 217, lifetime 108, 218, and transport behavior 219 can also be changed. Such phonon-exciton coupling has been widely studied in quantum dots and quantum wells 220, 221, 222, and more recently, in complex bulk and atomically-thin materials 215, 216, 217. By comparing the exciton lineshape in broadband absorption or photoluminescence spectra before and after the photoexcitation event, one can observe how excitons respond to phonons in the time domain, yielding insights for tuning exciton properties via ultrafast phonon engineering.
@@
For example, in anatase TiO2, a bulk semiconductor, Baldini et al. used photoinduced coherent acoustic phonons to control its excitons at room temperature, observing an exciton energy shift up to 30 meV, which is one of the largest reported in solids under a time-dependent perturbation 215. In a prototypical two-dimensional lead iodide perovskite, (PEA)2PbI4 (PEA = phenylethylammonium), it was further discovered that different excitons have varying coupling strengths with different coherent phonons, suggesting distinct polaronic characters of each exciton in this system and possibly in a broader class of hybrid semiconductors 216. In the atomically thin limit, Trovatello et al. also showed a pronounced phonon-exciton coupling effect in a monolayer MoS2, whose excitons are strongly affected by a photoinduced coherent phonon 217. Using transient absorption spectroscopy, the authors studied the interaction between out-of-plane phonon modes and excitons, noting a marked enhancement in the phonon oscillation amplitude at resonant exciton frequencies.
@@
The aforementioned studies of phonon-exciton coupling, however, lack spatial information, which is important for learning the transport behavior of excitons under the influence of phonons. Recently, Peng and coworkers demonstrated microscopic imaging of directional transport of interlayer excitons in bilayer WSe2 with the assistance of surface acoustic waves 219, and their experimental scheme is summarized in Fig. 9a. In this geometry, excitons were optically excited by a continuous-wave laser on the left edge of the sample and probed at different locations via photoluminescence imaging, giving rise to a spatially-resolved mapping of exciton propagation. As shown in Fig. 9b, the excitons were launched near the position at 20 m, where the strongest photoluminescence intensity appears. In the absence of a surface acoustic wave, the exciton is largely localized (Fig. 9b). However, by launching a surface acoustic wave using an interdigital transducer, excitons can travel a long distance (m, Fig. 9c), which is much longer than their typical diffusion length. Despite the lack of temporal resolution in this measurement, one can imagine a spatiotemporal experimental scheme by using a pulsed laser to launch the excitons. Crucially, this experiment introduces a new way to control exciton transport via electrically-excited or photoinduced acoustic waves, a versatile approach that can be adapted to many other 2D semiconductors and heterostructures.
@@ IV.5 Other collective excitation couplings
Despite the special focus on phonons as one of the coupled modes in the preceding sections, the full combinatorics of collective excitation coupling has been applied to many other scenarios such as magnon-magnon 225, 224, 223, 226, 227, exciton-magnon 228, plasmon-polariton 229, 148 among other interactions. In this section, we discuss a few examples that showcase the diversity of collective excitation couplings and their role in determining the ultrafast dynamics of quantum materials.
@@
Similar to nonlinear phononics, which opens a new pathway for the coherent manipulation of the crystalline structure, magnon-magnon coupling provides an avenue for controlling the spin dynamics. Utilizing 2D THz spectroscopy, Blank and coworkers revealed the down-conversion coupling of two magnon modes in the canted antiferromagnet FeBO3 (ref. 225). In another canted antiferromagnet, ErFeO3, Zhang and coworkers also demonstrated a unidirectional up-conversion coupling between distinct magnon modes 224. The canted spins in these materials result from the competition between antiferromagnetic interaction and Dzyaloshinskii-Moriya interaction, yielding a small net magnetic moment . As a result, there are two primary cooperative motions of the sublattice spins, corresponding to two distinct magnon modes sketched in Fig. 10a: the quasi-ferromagnetic (qFM) mode where precesses around the -axis, and the quasi-antiferromagnetic (qAFM) mode where oscillates. Given the different contributions to from the two modes, they can be selectively excited by tuning the direction of the magnetic field in the incident THz pump pulse, and their dynamics can be detected in the radiated free-induction decay signals. For example, when the magnetic field of the emitted THz pulse is parallel to that of the excitation field (), qFM mode and qAFM mode are selectively excited and observed when and , respectively (upper panel in Fig. 10b). This observation is expected as qFM involves a precession in the - plane while qAFM involves an amplitude oscillation along . Surprisingly, when , qAFM mode is still observed with (lower panel in Fig. 10b), which is not expected if there were only linear interactions between the driving THz magnetic field and magnons. To elicit the nonlinear interactions, 2D THz spectroscopy measurements were employed, where the nonlinear spectrum reveals a unidirectional up-conversion process from qFM to qAFM. As shown in Fig. 10c, the unidirectionality is evidenced by a single peak at the qFM excitation frequency and qAFM detection frequency, and importantly, a peak is absent when the corresponding excitation and detection frequencies are interchanged. Besides magnon up-conversion that echoes the phonon up-conversion seen in SrTiO3 (Fig. 5c, ref. 187), the sum and difference frequencies between these two magnon modes have been demonstrated in a cousin rare-earth ferrite, YFeO3, which are featured by a plethora of off-diagonal peaks in the nonlinear 2D terahertz spectrum in Fig. 10d (ref. 223). Like nonlinear phononics, these nonlinear effects of magnons demonstrated in the time domain help usher in new ways to coherently manipulate the spin degree of freedom in magnetic materials.
@@
The analogy between phonons and magnons goes beyond anharmonic couplings. Like phonons, a magnon wavepacket can propagate in materials, which forms the bedrock for spintronic and magnonic applications. When coupled with other collective excitations, magnon propagation can be modified. For instance, for magnons that are coupled with acoustic phonons, theoretical 230 and experimental 231, 232 works have shown that magnon polarons travel over long distances with a faster group velocity than magnons alone. In a study of a van der Waals antiferromagnetic semiconductor CrSBr, magnons were observed to coherently travel beyond several micrometers with a very fast group velocity, an effect that Bae and coworkers interpreted as the coupling between the magnon and acoustic phonon branches 228, 233. What is interesting about this experiment is that coherent magnons were probed by laser pulses tuned to the exciton resonances in CrSBr, demonstrating efficient optical access to spintronic information due to strong exciton-magnon couplings in this material. A subsequent experiment suggested that optically launched magnon propagation in CrSBr and likely other van der Waals magnets are instead mediated by long-range magnetic-dipole interactions 234, calling for additional spatiotemporally-resolved studies to understand magnon dynamics in this important class of materials.
@@
Non-local propagation effects mediated by the coupling between collective excitations are not restricted to magnons. Other examples include exciton propagation with the aid of surface acoustic waves in a GaAs quantum well 235, 236 or in a transition metal dichalcogenide 219, and surface-plasmon-polariton propagation at a metal-dielectric interface 237. Similar to a phonon-polariton, a plasmon-polariton is a hybridized excitation between charge oscillations and light. More specifically, when light illuminates a metal-dielectric interface, the oscillating electric field of light drives a collective oscillation of the surface charge, which in turn radiates electromagnetic waves. As a result, a hybridized wave, namely surface-plasmon-polariton that involves both the charge motion on the metal surface and electromagnetic field in the dielectric, propagates along the interface. To observe and control its propagation dynamics, PINEM is one of the leading techniques 150, 148. Figure 11a shows the spatiotemporally-resolved imaging of surface-plasmon-polaritons induced by near-infrared laser pulses on an isolated silver nanowire suspended on a graphene film 148, where surface-plasmon-polaritons are evidenced by the additional peaks that appear in the electron energy loss spectra when the light and electron pulses are temporally overlapped. These peaks result from the quantized gain and loss of energy when electrons interact with the surface-plasmon-polariton field. Using energy-filtered electrons, the temporal dynamics of the surface-plasmon-polariton field was imaged with high spatial resolution as it formed an evolving interferometric standing wave pattern (Fig. 11b–f). Moreover, with different light polarizations, different surface-plasmon-polariton field distributions were obtained, demonstrating the potential to control surface-plasmon-polaritons using light.
@@ V Summary and outlook
The study and manipulation of coupled collective excitations by ultrafast light pulses offer a new perspective on disentangling intertwined degrees of freedom that are difficult to isolate in equilibrium. In this review, we have highlighted several couplings that both yield an in-depth understanding of the equilibrium properties and enable the coherent control of the nonequilibrium dynamics in the host material.
While examples mentioned so far only include two participating collective excitations, multipartite couplings have been theoretically pondered, such as a hybrid exciton-phonon-photon that gives rise to a “phonoriton” in an optical cavity (Fig. 12a) 240, 238, and experimentally realized, such as the hybrid plasmon-exciton-photon — known as a “plexciton” — in WS2 (Fig. 12b), where cavity effects were achieved through nano-patterning 239, 241. These studies fall into a larger body of works that aims to manipulate quantum materials via cavity control, in which the light-matter coupling strength can be greatly enhanced by replacing classical laser fields with quantum-mechanical photon modes. One of the most celebrated examples is microcavity exciton–polaritons in GaAs quantum wells, which arise from the strong coupling between photons and excitons and exhibit remarkable effects such as Bose-Einstein condensation 242, 243, 244. More recently, a multitude of theoretical proposals have been put forward for tailoring the unique properties of quantum materials via cavity design 245, 246, with experimental progress in the study of topological states 247, superconductivity 248, and phase transitions 249 in different material platforms.
@@
First, advances in engineering the waveform of excitation light pulses — such as their polarization, carrier envelope phase, and orbital angular momentum — can greatly expand the toolkits for resonant excitation of collective modes and hence enable a deterministic control over material properties. A recent example is the synthesis of an intense, circularly-polarized THz pulse used to induce circular motions of atoms, otherwise known as chiral phonons, which can in turn induce a large effective magnetic field 76, 71, 70, 69. In particular, Davies et al. demonstrated that the induced field can significantly alter the magnetic order even at considerable distances from the phonon creation point 70, underscoring the need for deeper insight into chiral phonon transport. Equipped with the capability to deterministically excite particular phonon and magnon modes using the mechanisms introduced in this review, one can envision future experiments where chiral phonons and magnons are created at distinct sample locations, and their mutual interaction over space and time can be used to gain a mechanistic understanding of the propagation mechanism for realizing ultrafast non-local control of magnetic ordering.
@@
Second, from a technical point of view, great care needs to be taken when comparing the dynamics of different collective excitations probed by distinct experimental setups, making a quantitative study of their couplings challenging. An example of this issue arises when one extracts the electron-phonon coupling strength by comparing the electronic band change and atomic displacement, which are probed by time-resolved photoemission and X-ray diffraction respectively 19, 22. For the same incident pump fluence, the effective excitation fluence can differ substantially because photoemission is a surface-sensitive probe (sub- to few nanometers), whereas X-rays can probe much deeper into the bulk (tens of nanometers to micrometers). Therefore, careful fluence calibrations are needed when comparing these probes unless monolayer samples are used to eliminate the variance. An alternative route is to look for complementary probes with a similar probing depth, such as high-energy electron diffraction and extreme ultraviolet absorption spectroscopy, both operated under a transmission geometry. While the former is sensitive to the lattice degree of freedom, the latter excels in capturing carrier, magnetic, and excitonic excitations with exceptional time resolution 25, 250. One can envision a chamber design that integrates both probes in a single setup, making it possible to concurrently access lattice and electronic information following the same excitation pulse.
@@
Third, while the review has largely focused on coherent collective dynamics with a well-defined phase, not all light-induced collective excitations are coherent or fully reproducible over multiple pump pulses, particularly in systems with spatial inhomogeneities where defects and disorders are prevalent 251, 252, 253, 254. In this aspect, techniques with both spatial and temporal resolutions that operate in the single-shot mode are important for revealing the local, incoherent, and stochastic dynamics that vary from pulse to pulse, as recently demonstrated in the study of magnetic fluctuations in magnetic multilayer thin films 142, 143 and the heterogeneous insulator-to-metal transition in VO2 (ref. 144), both of which utilized a coherent X-ray source in large-scale user facilities. It is worth noting that stochastic dynamics can also be accessed without relying on single-shot detection 255, which makes it possible to study fluctuating systems with the stroboscopic scheme employed in the vast majority of ultrafast measurements. Along with rapid progress in facility-based ultrashort coherent X-ray sources, we also envision two key developments in table-top experiments for probing local and incoherent collective modes. At the extreme spatial scale, innovations in scanning tunneling microscopy have enabled new ways to image collective excitations at the atomic scale 145, 146, paving the way for controlling these local dynamics through creative engineering of nanoscopic defects. At the extreme temporal scale, ultrafast electron diffraction and microscopy, which have witnessed a rapid improvement of their temporal resolution 256, 109, 257, 258, 259, would provide a useful platform to study the initial emergence of incoherent phonons following photoexcitation 260, 261, which play an important role in determining the ensuing defect structure 53 and energy flow in the host material system 262.
@@
As we conclude this review that has touched a diverse set of experimental techniques and observables, it is apt to be reminded that the overarching goal of studying coupled collective excitations is to uncover a set of unifying principles that govern the interaction among the different degrees of freedom in solids. While studying collective excitations is certainly not the only way to unravel the intricate microscopic interactions, they offer us a way to describe a many-body system without the need to keep track of individual dynamics of an astronomical number of particles, making it tractable to study and engineer material properties both in and out of equilibrium.